OpenStudy (anonymous):

Solve, using the substitution method. y = 4x + 6 5x – y = 6 (Points : 1) There are an infinite number of solutions. There is no solution. (18, 12) (12, 54)

OpenStudy (luigi0210):

If the equations were the same they'd have infinitely many. But they're not so you can cross that out.

OpenStudy (luigi0210):

If they had the same slope they's be parallel and have no solution, so you can cross that one out too. Leaving just C and D as options.

OpenStudy (anonymous):

the answer tht i believe it is D because i dont think it could be lower then but im not entirely sure

OpenStudy (luigi0210):

Yup, that would be it~

OpenStudy (luigi0210):

Sub in: \[\LARGE 5x-(4x+6)=6\] Distribute: \[\LARGE 5x-4x-6=6\] Now just finish for x :3

OpenStudy (whpalmer4):

You could try actually solving it via the substitution method :-) \[y = 4x+6\]\[5x-y=6\]Substitute the first equation into the second:\[5x-(4x+6) = 6\]Collect like terms\[x-6=6\]\[x=12\]Now plug value of \(x=12\) into either equation to find \(y\) \[y=4(12)+6\]\[y=\]

OpenStudy (luigi0210):

If you want a good explanation^